# Generalizability of Graph Neural Network Force Fields for Predicting Solid-State Properties

### v1

@@ 1 Introduction
Machine learning approaches to aid molecular dynamics (MD) simulations have developed rapidly in recent years. Techniques such as deep learning have been applied to accelerated MD simulations [1, 2, 3, 4] to address their inherent time scale limitations. Creating accurate force fields has been a key area of development using deep learning architectures, often referred to as machine-learned force fields (MLFFs) or interatomic potentials (MLIPs). These models aim to achieve the scalability of classical MD simulations with the chemical accuracy of ab initio methods. Graph neural networks (GNNs) are a common architecture choice for MLFFs to encode the local atomic environments. Starting from SchNet [5], developments along this direction subsequently lead to DeepMD [1, 6, 7]. The ability to differentiate between charged and uncharged species, needed for modeling ionic solvents, was introduced by ANI [8, 9] and further developed into SANI [10, 11] and QRNN [12] to capture complex behavior such as polarization and chemical reactivity [13]. These MLFFs and MLIPs have been used to study a number of systems such as metals [14, 15, 16], biomolecules [17], polymers [18] and organic molecules [19]. Our recent work presented a systematic approach for testing the transferability of these MLFF/MLIP models through the TB-MLFF library [20]. However, existing tests primarily focus on liquid-phase structures and dynamics. Our prior solid-phase tests were limited to the phonon density of states (PDOS) for a perfect crystal at zero temperature. Since a significant potential application of MLFF/MLIP models involves finite-temperature modeling of solid materials with defect-controlled properties, more rigorous testing in this domain is essential before trusting these models for predictive computational materials science applications.
@@
This work introduces a more comprehensive set of solid-phase validation tests for MLFFs. We continue to use the GNN framework developed by Li et al. [21] to train an MLFF on data from the Lennard-Jones potential model of Ar, as a representative example. We extend our investigation beyond perfect crystals at zero temperature to include lattice vibration at finite temperatures and vacancy migration in imperfect crystals. On the finite-temperature lattice vibrations, we calculate the phonon dispersion relation, sensitive to both harmonic and anharmonic effects [22, 23], using the spectral energy density method (SED) [24, 25, 26] implemented in our C-XPCS library [27]. We also use the spectral energy density to estimate the thermal conductivity of the crystal [28] and benchmark the performance of the MLFF. On the defect dynamics in imperfect crystals, we calculate the vacancy jump rate from MD simulations using both the MLFF and the original potential model. The results are also compared against rate theory predictions  [29] with activation energy barrier computed from a modified version of the string method [30, 31, 32, 33, 34]. These tests allow us to benchmark the MLFF in the solid phase and can help determine the modifications to the dataset that might be required to make the MLFF more generalizable.
@@ 2.2 Spectral energy density
where $\tau_{0}$ is the total simulation time, and $N_{\rm T}$ is the total number of primitive cells in the simulation cell. $N_{\rm T}=N_{x}\times N_{y}\times N_{z}$, where $N_{x}$, $N_{y}$, and $N_{z}$ refer to the number of primitive cells in the $x$, $y$, and $z$ directions, respectively. $n_{i,j,k}$ denotes primitive cells with $i\in[1,N_{x}]$, $j\in[1,N_{y}]$, and $k\in[1,N_{z}]$. Eq. (7) is only applicable to crystals with a single-atom basis, such as the face-centered cubic (FCC) crystal of Ar considered here. $\dot{u}_{a}$ stands for the atomic velocity in the $a$-th direction, where $a=x$, $y$, or $z$. $\bm{r}_{0}(n_{i,j,k})$ denotes the equilibrium position of the $n_{i,j,k}$-th primitive cell. The resulting $\phi(\bm{k},\omega)$ map exhibits peaks at the ($\bm{k}$, $\omega$) loci that correspond to the phonon dispersion relation.
@@ 2.3 Vacancy Jump Rates
where $\nu_{0}$ is the rate prefactor and $E_{\rm b}$ is the energy barrier. There are various ways to estimate the rate prefactor, such as the Eyring-Polanyi equation [29] and the harmonic transition state theory [36]. A rough estimate for the rate prefactor is the Debye frequency $\nu_{\rm D}$, which is related to temperature $\Theta_{\rm D}$ by $\nu_{\rm D}=k_{\rm B}\Theta_{\rm D}/h$, where $h=2\pi\hbar$ is Planck’s constant. Given that the Debye temperature for solid Argon is $\Theta_{\rm D}=92$ K [37], the Debye frequency is $\nu_{\rm D}=1.92\times 10^{12}$ s-1. The energy barrier is the energy increase from the initial state (A) to the transition state (S), which can be separately predicted by minimum energy path (MEP) searches, such as the nudged-elastic band (NEB) method [38, 39]. These estimates of $\nu_{0}$ and $E_{\rm b}$ serve as additional benchmarks to the MLFF’s ability to accurately capture vacancy jump rates.
@@ 3 Results
We present comparisons between the GNN-based MLFF and the original Lennard-Jones (LJ) potential on simulations of solid Argon, which forms a face-centered cubic (FCC) lattice at low temperatures. The GNN is trained on the LJ force field using samples containing both solid and liquid configurations (from 10 K to 105 K), similar to the previous models [27]. The solid configurations used in the training of GNN-based MLFF correspond to perfect crystals (i.e. vacancy-free) unless mentioned otherwise explicitly.
@@ 3.1 Phonon dispersion relations
As discussed earlier in Secton 2.1, the Hessian matrix method can be used to obtain the phonon density of states at 0 K. Using a finite difference step size of $\Delta x=3.4\times 10^{-4}\,$  for the GNN-based MLFF, we obtain a nearly symmetric $\mathcal{H}$ (the deviation of symmetricity $\mathcal{S}$ from $1$ is on the order of $10^{-5}$). The coarser finite difference step size $\Delta x$ doesn’t affect the agreement of the PDOS from the GNN-based MLFF against the reference LJ potential, as shown in Fig. 1. The mean absolute error (MAE) for each individual eigenfrequency (for the same eigenmode) is under 0.03 THz and the maximum absolute error (MaxAE) is less than 0.1 THz.
@@
The Hessian matrix allows us to calculate the phonon dispersion at 0 K where the phonons are trapped in a given mode of vibration (infinite lifetime). To obtain the phonon dispersion relation at a finite temperature where the phonons have a finite lifetime in any particular mode, we use the spectral energy density (SED) method (see Section 2.2). To this end, we consider a supercell that contains $5\times 5\times 100$ FCC unit cells, with a lattice parameter, $a=4.808$. The simulation cell is extended significantly in the $z$-direction to achieve a fine resolution in $k_{z}$. This quasi-1D configuration allows us to compute the phonon dispersion as a function of $k_{z}\in[-\pi/a,\pi/a)$ (from $\Gamma$ point to X point in the reciprocal lattice). The initial configuration is relaxed to an energy minimum configuration by using the conjugate gradient algorithm, to obtain the equilibrium position, $\bm{r}(n_{x,y,z},0)$, for each primitive cell. We use the equilibrium position calculated by the LJ force field for the SED calculations for both the MD simulation and GNN-MD simulation trajectory. We first perform an NVT simulation thermostatted at 40 K (finite temperature) and adjust the simulation cell dimensions every 10 ps to reach the target pressure of $p=137.1$ bar. We then obtain the equilibrium lattice constant of the crystal at finite temperature by performing energy relaxation. We finally perform an NVT simulation at the equilibrium box size for 100 ps, thermostatted at $T=40$ K. The equilibrated configuration is used as the initial configuration for the MD and GNN-MD simulations, each carried out under an NVT ensemble thermostatted at 40 K for 107.8 ps, with configuration snapshots being recorded at 107.8 fs intervals. We use these 1000 frames to compute the phonon dispersion relation using the SED analysis, which is sufficient since it is greater than the phonon lifetimes of solid Argon at this temperature.
@@ 3.2 Thermal conductivity predictions
The SED signal generated from the MD and GNN-MD trajectory can be used to extract the thermal conductivity, $\kappa$, of the system using the relation in Eq. (9). We use a cubic simulation cell for the $\kappa$ calculations, where the supercell is made of $10\times 10\times 10$ FCC unit cells and is thermostatted at 40 K and barostatted at a pressure of 137.1 bar. We limit the size of the simulation cell in the interest of the computational expense of the GNN-MD simulation. However, our results confirm that the simulation cell is large enough for a reliable estimate of $\kappa$. To validate the $\kappa$ obtained from the SED calculations, we compare it against reference calculations from an MD simulation of a larger simulation cell, using the Green-Kubo formulation [40]. The reference simulation cell consists of $50\times 50\times 50$ FCC unit cells thermostatted at 40 K and barostatted at a pressure of 137.1 bar and the resulting thermal conductivity is $\kappa\sim 0.49$ W m-1K-1. From the SED calculations, we obtain $\kappa_{\rm MD}\sim 0.45$ W m-1K-1 from the MD trajectory and $\kappa_{\textrm{GNN-MD}}\sim 0.55$ W m-1K-1 from the GNN-MD trajectory, which are both within $15\%$ of the reference calculations. Overall, the phonon dispersion and thermal conductivity calculations from the MD and GNN-MD trajectories are in good agreement. These results provide an estimate of the quality of the GNN-based MLFF in thermal conductivity predictions and illustrate the value of these benchmarking tests for evaluating solid-state behavior in MLFF or MLIP models.
@@ 3.3 Vacancy jump rates
Having tested the performance of the MLFF for perfect crystals, we now present the analysis of the vacancy jumping rate in an imperfect crystal (containing a single vacancy). We started with a 256-atom solid configuration (in FCC lattice) and then removed one atom to create a vacancy. We also imposed a negative pressure of $\sim 2.5$ kbar, resulting in a low-density simulation cell at $5.711$ mol/cm3. The low-density simulation cell is chosen to reduce the energy barrier for vacancy jumps, so that they can be observed in the limited timescales accessible by MD and GNN-MD simulations. We use the resulting configuration to run MD and GNN-MD simulations for $\sim 53.9$ ns where we record the configuration snapshot every $\sim 1.078$ ps. We employ the Wigner-Seitz analysis using OVITO [41] to identify the vacancy in each configuration by comparing it to a perfect crystal configuration. A vacancy jump is recorded whenever the atom ID of the zero-occupancy atom changes in the Wigner-Seitz analysis.
@@ 4 Conclusions
The tests presented here demonstrate the promise of MLFFs for studying defects in solids and rare events, which are important for applications in material science. These tests can be used to identify the modifications to the dataset that might be required to make the MLFF more generalizable, making high throughput and high-fidelity development of MLFF possible.
@@ Acknowledgements
S.M. and W.C. acknowledge support from the Precourt Pioneering Project of Stanford University. S.M. and W.C. acknowledge support from the Air Force Office of Scientific Research under award number FA9550-20-1-0397. We would like to thank Klara Suchan for the discussions on the spectral energy density method.

### v2

@@ 1 Introduction
Machine learning approaches to aid molecular dynamics (MD) simulations have developed rapidly in recent years. Techniques such as deep learning have been applied to accelerated MD simulations [1, 2, 3, 4] to address their inherent time scale limitations. Creating accurate force fields has been a key area of development using deep learning architectures, often referred to as machine-learned force fields (MLFFs) or interatomic potentials (MLIPs). These models aim to achieve the scalability of classical MD simulations with the chemical accuracy of ab initio methods. However, verifying that the machine-learned models trained on data from simpler and less expensive empirical potentials can reproduce their behaviors can lead to valuable insights and important guidelines for constructing reliable MLFFs and MLIPs [5]. Graph neural networks (GNNs) are a common architecture choice for MLFFs to encode the local atomic environments (short and long-range interactions). Starting from SchNet [6], developments along this direction subsequently lead to DeeP-MD [1, 7, 8]. The ability to differentiate between charged and uncharged species, needed for modeling ionic solvents, was introduced by ANI [9, 10] and further developed into SANI [11, 12] and QRNN [13, 14] to capture complex behavior such as polarization and chemical reactivity [15]. These MLFFs and MLIPs have been used to study a number of systems such as metals [16, 17, 18], biomolecules [19], polymers [14] and organic molecules [20]. Our recent work presented a systematic approach for testing the transferability of these MLFF/MLIP models through the TB-MLFF library [5]. However, existing tests primarily focus on liquid-phase structures and dynamics. Our prior solid-phase tests were limited to the phonon density of states (PDOS) for a perfect crystal at zero temperature. Since a significant potential application of MLFF/MLIP models involves finite-temperature modeling of solid materials with defect-controlled properties, more rigorous testing in this domain is essential before trusting these models for computational materials science applications.
@@
This work introduces a more comprehensive set of solid-phase validation tests for MLFFs. We continue to use the GNN framework developed by Li et al. [21] to train an MLFF on data from the Lennard-Jones potential model of Ar, as a representative example. We extend our investigation beyond perfect crystals at zero temperature to include lattice vibration at finite temperatures and vacancy migration in imperfect crystals. On the finite-temperature lattice vibrations, we calculate the phonon dispersion relation, sensitive to both harmonic and anharmonic effects [22, 23], using the spectral energy density method (SED) [24, 25, 26] implemented in our C-XPCS library [27]. Similar calculations have also recently been implemented using DeeP-MD simulations [28, 29]. We also use the spectral energy density to estimate the thermal conductivity of the crystal [30] and benchmark the performance of the MLFF. Regarding defect dynamics and the analysis of rare events in imperfect crystals, we calculate the vacancy jump rate from MD simulations using both the MLFF and the original potential model. The results are also compared against rate theory predictions  [31] with activation energy barrier computed from a modified version of the string method [32, 33, 34, 35, 36]. These tests allow us to benchmark the MLFF in the solid phase and can help determine the modifications to the dataset that might be required to make the MLFF more generalizable.
While in classical force fields, the atomic forces are computed as the negative derivatives of a potential energy function, some of the GNN-based models directly compute the forces without having a potential energy [21]. In other words, the forces by these GNN-based models may be non-conservative, i.e. a potential energy function may not exist. In this case, one may argue whether such models can be designated as force fields (MLFFs), or perhaps calling them “force predictors” is more accurate. In this paper, we shall still refer to such models as MLFFs, in keeping with the existing literature. Furthermore, since such models are trained on data obtained from conservative forces, their predicted forces are nearly (although not exactly) conservative, so that the forces should be close to the negative derivatives of some potential energy function.
@@ 2.2 Spectral energy density
where $\tau_{0}$ is the total simulation time, $m$ is the atomic mass, and $N_{\rm T}$ is the total number of primitive cells in the simulation cell. $N_{\rm T}=N_{x}\times N_{y}\times N_{z}$, where $N_{x}$, $N_{y}$, and $N_{z}$ refer to the number of primitive cells in the $x$, $y$, and $z$ directions, respectively. $n_{i,j,k}$ denotes primitive cells with $i\in[1,N_{x}]$, $j\in[1,N_{y}]$, and $k\in[1,N_{z}]$. Eq. (7) is only applicable to crystals with a single-atom basis, such as the face-centered cubic (FCC) crystal of Ar considered here. $\dot{u}_{a}$ stands for the atomic velocity in the $a$-th direction, where $a=x$, $y$, or $z$. $\bm{r}_{0}(n_{i,j,k})$ denotes the equilibrium position of the $n_{i,j,k}$-th primitive cell. The resulting $\phi(\bm{k},\omega)$ map exhibits peaks at the ($\bm{k}$, $\omega$) loci that correspond to the phonon dispersion relation.
@@ 2.3 Vacancy Jump Rates
where $\nu_{0}$ is the rate prefactor and $E_{\rm b}$ is the energy barrier. There are various ways to estimate the rate prefactor, such as the Eyring-Polanyi equation [31] and the harmonic transition state theory [38]. A rough estimate for the rate prefactor is the Debye frequency $\nu_{\rm D}$, which is related to temperature $\Theta_{\rm D}$ by $\nu_{\rm D}=k_{\rm B}\Theta_{\rm D}/h$, where $h=2\pi\hbar$ is Planck’s constant. Given that the Debye temperature for solid Argon is $\Theta_{\rm D}=92$ K [39], the Debye frequency is $\nu_{\rm D}=1.92\times 10^{12}$ s-1. The energy barrier is the energy increase from the initial state (A) to the transition state (S), which can be separately predicted by minimum energy path (MEP) searches, such as the string method [32, 33, 34, 35, 36]. These estimates of $\nu_{0}$ and $E_{\rm b}$ serve as additional benchmarks to the MLFF’s ability to accurately capture vacancy jump rates.
@@ 3 Results
We present comparisons between the GNN-based MLFF and the original Lennard-Jones (LJ) potential on simulations of solid Argon, which forms a face-centered cubic (FCC) lattice at low temperatures. The GNN is trained on the LJ force field using samples containing both solid and liquid configurations (from 10 K to 105 K), similar to the previous models [27]. The solid configurations used in the training of GNN-based MLFF correspond to perfect crystals (i.e. vacancy-free) unless mentioned otherwise explicitly. The MD simulations are performed in non-dimensional or LJ units where the units in distance, time, and energy are given by: $3.41\sigma=$ 3.4 Å, $2.1561\tau=$ 2.156 ps, and $0.0103271\epsilon=$ 0.010327 eV, respectively.
@@ 3.1 Phonon dispersion relations
As discussed earlier in Secton 2.1, the Hessian matrix method can be used to obtain the phonon density of states at 0 K. Using a finite difference step size of $\Delta x=3.4\times 10^{-4}\,$  ($10^{-4}\sigma$) for the GNN-based MLFF, we obtain a nearly symmetric $\mathcal{H}$ (the deviation of symmetricity $\mathcal{S}$ from $1$ is on the order of $10^{-5}$). This result indicates that the forces predicted by the MLFF are nearly (although not exactly) conservative. The coarser finite difference step size $\Delta x$ doesn’t affect the agreement of the PDOS from the GNN-based MLFF against the reference LJ potential, as shown in Fig. 1. The mean absolute error (MAE) for each individual eigenfrequency (for the same eigenmode) is under 0.03 THz and the maximum absolute error (MaxAE) is less than 0.1 THz.
@@
The Hessian matrix allows us to calculate the phonon dispersion at 0 K where the phonons are trapped in a given mode of vibration (infinite lifetime). To obtain the phonon dispersion relation at a finite temperature where the phonons have a finite lifetime in any particular mode, we use the spectral energy density (SED) method (see Section 2.2). To this end, we consider a supercell that contains $5\times 5\times 100$ FCC unit cells, with a lattice parameter, $a=4.808$($\sqrt{2}\sigma$) . The simulation cell is extended significantly in the $z$-direction to achieve a fine resolution in $k_{z}$. This quasi-1D configuration allows us to compute the phonon dispersion as a function of $k_{z}\in[-\pi/a,\pi/a)$ (from $\Gamma$ point to X point in the reciprocal lattice). The initial configuration is relaxed to an energy minimum configuration by using the conjugate gradient algorithm, to obtain the equilibrium position, $\bm{r}(n_{x,y,z},0)$, for each primitive cell. We use the equilibrium position calculated by the LJ force field for the SED calculations for both the MD simulation and GNN-MD simulation trajectory. We first perform an NVT simulation thermostatted at 40 K (finite temperature) and adjust the simulation cell dimensions every 10 ps to reach the target pressure of $p=137.1$ bar. We then obtain the equilibrium lattice constant of the crystal at finite temperature by performing energy relaxation. We finally perform an NVT simulation at the equilibrium box size for 100 ps, thermostatted at $T=40$ K. The equilibrated configuration is used as the initial configuration for the MD and GNN-MD simulations, each carried out under an NVT ensemble thermostatted at 40 K for 107.8 ps ($50\tau$), with configuration snapshots being recorded at 107.8 fs ($0.05\tau$) intervals. We use these 1000 frames to compute the phonon dispersion relation using the SED analysis, which is sufficient since it is greater than the phonon lifetimes of solid Argon at this temperature.
@@ 3.2 Thermal conductivity predictions
The SED signal generated from the MD and GNN-MD trajectory can be used to extract the thermal conductivity, $\kappa$, of the system using the relation in Eq. (9). We use a cubic simulation cell for the $\kappa$ calculations, where the supercell is made of $10\times 10\times 10$ FCC unit cells and is thermostatted at 40 K and barostatted at a pressure of 137.1 bar. We limit the size of the simulation cell in the interest of the computational expense of the GNN-MD simulation. However, our results confirm that the simulation cell is large enough for a reliable estimate of $\kappa$. To validate the $\kappa$ obtained from the SED calculations, we compare it against reference calculations from an MD simulation of a larger simulation cell, using the Green-Kubo formulation [40]. The reference simulation cell consists of $50\times 50\times 50$ FCC unit cells thermostatted at 40 K and barostatted at a pressure of 137.1 bar and the resulting thermal conductivity is $\kappa\sim 0.49$ W m-1K-1. From the SED calculations, we obtain $\kappa_{\rm MD}\sim 0.45$ W m-1K-1 from the MD trajectory and $\kappa_{\textrm{GNN-MD}}\sim 0.55$ W m-1K-1 from the GNN-MD trajectory, which are both within $15\%$ of the reference calculations. Note that the thermal conductivity calculations are carried over all the $\phi(\bm{k},\omega)$ pairs on the 3D Fourier grid (wave-vector space) and are not just restricted to the high symmetry directions. Overall, the phonon dispersion and thermal conductivity calculations from the MD and GNN-MD trajectories are in good agreement. These results provide an estimate of the quality of the GNN-based MLFF in thermal conductivity predictions and illustrate the value of these benchmarking tests for evaluating solid-state behavior in MLFF or MLIP models.
@@ 3.3 Vacancy jump rates
Having tested the performance of the MLFF for perfect crystals, we now present the analysis of the vacancy jumping rate in an imperfect crystal (containing a single vacancy). We started with a 256-atom solid configuration (in FCC lattice) and then removed one atom to create a vacancy. We also imposed a negative pressure of $\sim 2.5$ kbar, resulting in a low-density simulation cell at $5.711$ mol/cm3. The low-density simulation cell is chosen to reduce the energy barrier for vacancy jumps, so that they can be observed in the limited timescales accessible by MD and GNN-MD simulations. We use the resulting configuration to run MD and GNN-MD simulations for $\sim 53.9$ ns ($2.5\times 10^{4}\,\tau$) where we record the configuration snapshot every $\sim 1.078$ ps ($0.5\,\tau$). We employ the Wigner-Seitz analysis using OVITO [41] to identify the vacancy in each configuration by comparing it to a perfect crystal configuration. A vacancy jump is recorded whenever the atom ID of the zero-occupancy atom changes in the Wigner-Seitz analysis.
@@ 4 Conclusions
In future work, addressing the challenges posed by the limitations of GNN-based force fields requires a careful analysis of data coverage within the configuration space. This work suggests that the issue lies not in data size but rather in ensuring sufficient diversity to cover a wide range of configurations. For example, the current model explored here is likely to struggle with scenarios such as free surfaces, where the underlying potential energy landscape differs significantly from the training data, and other MLFF architectures may be worth exploring [42]. Extending this analysis to MLFFs fitted to data generated by more complex potentials, such as the Tersoff potential or EAM potential, could provide additional insights, though we anticipate increased difficulties in accurately capturing many-body effects. Finally, developing robust error quantification metrics depending on physical observables (such as thermal conductivity and vacancy jump rates) beyond forces and energies will be crucial to systematically assess model performance and identify sources of error. Exploring these directions will help guide improvements in the design and training of GNN-based force fields or interatomic potentials. The tests presented here demonstrate the promise of MLFFs for studying defects in solids and rare events, which are important for applications in material science. Although the GNN-based model benchmarked here is trained on force data from the simple LJ potential, the workflow of training and testing adopted here can be equally applied to MLFFs trained on any force data, including atomic forces obtained from ab initio models. These tests can be used to identify the modifications to the dataset that might be required to make the MLFF more generalizable, making high throughput and high-fidelity development of MLFF possible.
@@ Acknowledgements
S.M. and W.C. acknowledge support from the Precourt Pioneering Project of Stanford University. S.M. and W.C. acknowledge support from the Air Force Office of Scientific Research under award number FA9550-20-1-0397. We would like to thank Klara Suchan and Myung Chul Kim for their discussions on the manuscript.